By changing the independent variable.solve the differential equation. d2x/dx2 - (dy/dx)(1/x) + 4x2y = x4

1
Expert's answer
2020-07-12T18:10:26-0400

Let "t=x^2,x=\sqrt{t}"

"y''-y'/(\sqrt{t})+4yt=t^2"

"\frac{dy}{dx}=2\sqrt{t}\frac{dy}{dt},\frac{d^2y}{dx^2}=4t\frac{d^2y}{dt^2}+2\frac{dy}{dt}"


"4t\frac{d^2y}{dt^2}+4ty(t)=t^2"

"\frac{d^2y}{dt^2}+y(t)=t/4"


"k^2+1=0"

"k=\pm i"

The general solution:

"y(t)=c_1cost+c_2sint"

For the particular solution:

"\tilde{y}(t)=A+Bt"

Then:

"A+Bt=t\/4"

"A=0,B=1\/4"

"\tilde{y}(t)=t/4"


So:

"y(t)=t\/4+c_1cost+c_2sint"


Answer:

"y(x)=x^2\/4+c_1cos(x^2)+c_2sin(x^2)"